117. Find the equation for each of the graphs below: 1 4 A B.
Answer #1

A) we have to write the equation of given graph

As we can see from the graph that the function is repeate itself after period of 4π.

Then w= 2π/T = 2π/4π = 1/2

It has the maximum amplitude of (A) = 3

Zero shifting.

Now equation of the curve

y = A*cos(wt) = 3cos(t/2)

B) Given graph repeat itself after period of π so

T= π and w= 2π/T = 2π/π = 2

As given graph is inverted graph of sine function

A= 4

Now equation of curve

y = -A*sin(wt) = -4*sin(2t)